OpenStudy (anonymous):

What 2 number multiply to get 4440 but add to get 9?

OpenStudy (lgbasallote):

xy = 4440 x + y = 9 from x + y = 9 we can say that x = 9 - y substitute that in the first equation... \[(9-y)y = 4440\] did that help?

OpenStudy (anonymous):

I just need the answer really quick

OpenStudy (anonymous):

he gave you as close of an answer as he can give you... We're not here to help you cheat your way through online classes. We are here to interact and teach

OpenStudy (anonymous):

I know. I am just not good with math and when i solved that I got y=-4440/y -9 and I dont think that's correct